# DefendCLI: {Command-Line} Driven Attack Provenance Examination

### v1

@@ Abstract
Endpoint Detection and Response (EDR) solutions embrace the method of attack provenance graph to discover unknown threats through system event correlation. However, this method still faces some unsolved problems in the fields of interoperability, reliability, flexibility, and practicability to deliver actionable results. Our research highlights the limitations of current solutions in detecting obfuscation, correlating attacks, identifying low-frequency events, and ensuring robust context awareness in relation to command-line activities. To address these challenges, we introduce DefendCLI, an innovative system leveraging provenance graphs that, for the first time, delves into command-line-level detection. By offering finer detection granularity, it addresses a gap in modern EDR systems that has been overlooked in previous research. Our solution improves the precision of the information representation by evaluating differentiation across three levels: unusual system process calls, suspicious command-line executions, and infrequent external network connections. This multi-level approach enables EDR systems to be more reliable in complex and dynamic environments. Our evaluation demonstrates that DefendCLI improves precision by approximately $1.6\times$ compared to the state-of-the-art methods on the DARPA Engagement Series attack datasets. Extensive real-time industrial testing across various attack scenarios further validates its practical effectiveness. The results indicate that DefendCLI not only detects previously unknown attack instances, which are missed by other modern commercial solutions, but also achieves a $2.3\times$ improvement in precision over the state-of-the-art research work.
Index Terms:
Endpoint Detection and Response, Attack Investigation, Advanced Persistent Threat.
@@ I Introduction
To provide a more automated security orchestration scheme, we integrate Large Language Models (LLMs) into our threat recommendation system to generate rapid preliminary assessments, enabling efficient triage before comprehensive threat analysis. Extensive experiments and evaluations show that DEFENDCLI effectively manages real-world cyber threats by identifying previously undetected attack instances. Our contributions are summarized as follows:
@@ II Related Work
The attack provenance graph is a powerful and comprehensive tool that maps the origins, evolution, and interconnections of entities and events within a system, with a particular emphasis on data flow and processes [6, 7]. By constructing detailed network diagrams, it traces the journey of data, illustrating how information is accessed, transmitted, and modified across various endpoints.
We argue that, despite significant progress made by existing methods in leveraging provenance analysis for system monitoring, their effectiveness in detecting and explaining command-line activities remains limited. These activities, however, often provide actionable insights for threat confirmation, which are closely related to the RQs discussed above.
This section provides a comparative analysis of related work by discussing their limitations across four key dimensions [8] that significantly impact the effectiveness of the command-line activity detection: restricted effectiveness obfuscation detection (OD), unclear and insufficient attack correlation (AC), incapable of low-frequency events detection (LFD), and limited context awareness (CA).
Table I qualitatively compares DefendCLI with previous methods in their granularity levels to detect command-line activities, based on their respective algorithmic designs (AD). A more rigorous quantitative evaluation is provided in Section V for the evidential validation.
TABLE I: Comparison of DEFENDCLI and Other Methods in Command-Line Detection (Granularity Levels). : Dominant, : Decent, : Moderate, ❍: Basic.
@@
1
OD: Restricted Effectiveness in Obfuscation Detection. [15] A significant limitation of many previous approaches lies in their heavy reliance on detecting structural differences or anomalies within execution graphs to identify malicious activities. Although this strategy can be effective for well-defined and previously observed attack patterns, it falls short when addressing the challenges posed by obfuscated command-line execution.
Attackers often leverage obfuscation techniques to disguise malicious behavior, making it difficult for detection systems to identify these threats using traditional anomaly-based or structural methods. The lack of robustness against obfuscation represents a critical vulnerability in these systems, which attackers can exploit to evade detection.
Detecting obfuscated commands is essential for mitigating advanced attack techniques that often rely on subtle modifications to otherwise legitimate processes. However, most existing methods struggle to achieve satisfactory results in this area. For example, Holmes [9] demonstrates decent effectiveness in identifying some obfuscation patterns, but its scope remains limited and fails to provide a comprehensive solution due to its rule-based nature. On the other hand, Unicorn [10], DepComm [11], ProGrapher [12], Nodlink [13], and DISTDET [14] show basic performance in this domain. These methods primarily focus on detecting more straightforward anomalies or structural changes and are not designed to handle sophisticated obfuscation techniques, such as polymorphic commands, encoded payloads, or the use of benign system utilities for malicious purposes. This limitation significantly reduces their effectiveness in combating modern, highly evasive attack strategies, leaving critical gaps in system security.
2
AC: Unclear and Insufficient Attack Correlation. [16] Similarly, these methods often struggle with attack correlation, as identifying relationships between commands typically requires a contextual understanding that goes beyond detecting mere structural deviations. Effective attack correlation is crucial for piecing together the sequence of events within an attack, enabling security teams to understand the full scope and intent of malicious activities.
Reconstructing attack chains through command correlation is vital for analyzing complex threat patterns and attributing malicious activities to specific campaigns. Among existing approaches, Nodlink [13] demonstrates dominant performance by using a steiner tree framework to correlate node-level events. Holmes [9] also has decent success, employing attack chain reconstruction techniques that provide a cohesive view of adversarial behavior. Similarly, Unicorn [10], DepComm [11], and ProGrapher [12] employ graph-based or dependency-oriented methods to trace structural relationships and event dependencies. While effective to a degree, these methods often fall short in capturing the nuanced contextual information necessary to identify more sophisticated and stealthy attack patterns.
In contrast, DISTDET [14] performs basically in this domain due to its lack of robust correlation mechanisms, as it simply deduplicates similar command-line entries without distinguishing between malicious and benign instances. Its limited ability to interpret complex relationships between commands diminishes its effectiveness in reconstructing multi-stage attack chains. This shortcoming is particularly problematic for detecting advanced persistent threats (APTs), which often rely on well-coordinated, multi-layered tactics to achieve their objectives. Overall, the inability of existing methods to fully capture and correlate command-line interactions represents a significant barrier to achieving comprehensive threat detection and analysis.
3
LFD: Incapable of Low-Frequency Events Detection. [17] Furthermore, detecting low-frequency patterns, which are often indicative of novel or highly targeted attacks, remains a critical gap in existing solutions. These patterns typically lack the historical context or recurrence that anomaly-based systems rely upon, making them harder to identify with conventional techniques. Methods that depend solely on anomaly detection tend to produce high false positive rates or miss subtle, context-dependent signals that require a deeper understanding of system behavior.
Detecting rare or low-frequency command-line patterns is particularly vital for identifying emerging threats and sophisticated targeted attacks. However, most existing approaches fail to effectively address this dimension. For instance, Holmes [9] offers moderate rule-based support for low-frequency event detection. Similarly, Unicorn [10], DepComm [11], ProGrapher [12], Nodlink [13], and DISTDET [14] primarily rely on identifying common patterns based on historical profiles of normal activity. However, they overlook the complexities of rare or context-specific anomalies, which frequently evolve and pose significant security challenges. This limitation hampers their ability to detect and respond to advanced stealth attacks that exploit the blind spots of conventional detection mechanisms.
4
CA: Limited Context Awareness. [18] Finally, the lack of context awareness in many existing approaches significantly undermines their ability to differentiate between benign and malicious commands, particularly in dynamic environments where identical commands can have vastly different implications depending on their execution context. Context awareness is critical for accurately interpreting command behavior, as the same command might serve legitimate purposes in one situation but be part of an attack in another. Without the ability to incorporate and analyze contextual data, detection systems often generate high false positive rates or fail to detect subtle malicious activities embedded within otherwise benign operations.
Context awareness enables detection systems to consider factors such as the command’s origin, associated processes, user privileges, and the broader system state, thereby reducing false alarms and enhancing detection precision. Among existing methods, Holmes [9] provides a moderate rule-based context awareness for known attacks, which limits its ability to distinguish between benign and unknown malicious command usage effectively. Similarly, Unicorn [10], ProGrapher [12], Nodlink [13], and DISTDET [14] exhibit basic performance in this area due to minimal integration of contextual data into their detection processes. These methods primarily rely on structural or anomaly-based techniques without considering the broader operational environment in which commands are executed.
In contrast, DepComm [11] shows decent performance using partial contextual information within their method. For instance, the method may take into account some system attributes as dependencies to reduce the attack irrelevance. However, its reliance on context reduction often fails to capture the full spectrum of environmental variables, making them inadequate for detecting sophisticated threats that exploit contextual subtleties. The lack of robust context-aware mechanisms in these approaches leaves significant blind spots in detecting stealthy tactics that mimic legitimate behavior, allowing attackers to exploit contextual ambiguities and bypass traditional detection methods.
III Detection Surface
Figure 3 presents the detection surface of our solution, highlighting its effectiveness compared to previous research, which struggled to achieve accurate detection. DefendCLI focuses on three levels of detection to thoroughly monitor unusual activities, interacting with system audit logs: (1) PC: it identifies unusual system process calls across the OS-level for each endpoint; (2) CL: it analyzes each individual command-line execution within the detected system process chain; (3) NC: it provides information on network connections related to each system process. However, DefendCLI is not a silver bullet to detect all malicious behaviors as part of the advanced security analysis function in EDR. It is often necessary to combine it with other security devices, such as Anti-Virus [19] and NIDS [20], to achieve more comprehensive results. Our solution focuses primarily on attack activities that can be observed and monitored from system audit logs, which provide detailed behavioral and operational information.
1
PC: Unusual Interacted System Process Calls. Observing system process calls is crucial for detecting abnormal OS activities, as legitimate software typically does not invoke unrelated applications or scripts. Despite attackers embedding malicious code in trusted software, a technique known as Living off the Land Attack [21], process chains still provide transparency by exposing deviations from normal behavior [22]. For instance, when malware is injected into a legitimate process, it spawns a new thread within the same namespace, altering system memory allocation. If a thread attempts to access another process’s memory, the system audit log records it as interprocess access, making such activities detectable through process call interactions.
Figure 3: Detection Surface: Given two similar attack chains, previous detection methods failed to distinguish the structural differences, leading to an inability to correctly identify the mimicry attack (right) resulted in blind classification.
2
CL: Suspicious Command-Line Executions. The effectiveness of system process chains is often compromised when malware embeds itself within legitimate processes, reducing visibility and detectability. Our solution addresses this by analyzing command-line attributes of flagged system processes within a process chain. This approach, which inspects command-line inputs from both interactive shells and OS API executions, enhances security analysts’ ability to swiftly review EDR results when process chain investigations become challenging. Certain command-line interactions are critical for OS functions and attack phases, making them unavoidable despite evasion tactics.
For example, the Turla APT Group [23] employs malware that leverages .NET or C++ to communicate with C2 servers, concealing malicious command-line activity. However, system process call analysis can still reveal anomalies, such as, Explorer.exe $\rightarrow$ Powershell $\rightarrow$ csc.exe (command-line: /noconfig…4krwc2ua.cmdline) $\rightarrow$ WMI.exe (command-line: tasklist). Notably, csc.exe rarely invokes 4krwc2ua.cmdline from the .NET library and should not trigger WMI.exe to execute tasklist, marking it as a potential anomaly.
3
NC: Rarely Seen External Network Connections. To boost detection capabilities, we integrate basic network data along with process-level and command-line-level analysis. Unlike network intrusion detection that uses detailed network attributes from packet inspection, the system audit log only records the essential network information IP: Port. The advantage of this method is its ability to link network connections directly with specific system processes. In the context of the Trula APT case, the detection outcome will be supplemented with network connections such as, Explorer.exe $\rightarrow$, Powershell $\rightarrow$, csc.exe (command-line: /noconfig/…//4krwc2ua.cmdline) $\rightarrow$, WMI.exe (command-line: tasklist) $\rightarrow$, and connection: 154.26.156[.]62:4444. This series reveals an unusual external connection by csc.exe, a fundamental component of .NET, to execute system commands through the 4krwc2ua.cmdline library object.
IV DefendCLI
To address the above challenges, we propose DefendCLI, a comprehensive system built with significant engineering effort to support actionable threat verification. We begin by outlining the system architecture, then describe each component following its execution workflow. For each part, we explain the design rationale and discuss the techniques used to implement our solution.
IV-A System Overview
@@
Figure 4 illustrates the architecture of the DefendCLI system, which adopts a highly modular design to facilitate extensibility and clarity. The system is composed of two core components: Attack-Clause Sketch, responsible for constructing the attack provenance graph, and Attack-Evidence Awareness, dedicated to threat detection.
1
Attack-Clause Sketch. The attack-clause sketch can be constructed to systematically represent and trace the logical structure of an attack as discussed below.
IV-A1 Attack Provenance Construction
It begins with the attack provenance construction module, which is designed to capture the interactive relationships among multi-level system events. This module is carried out in three key steps: (1) attack-irrelevance reduction: a reduction method is used to simplify the graph while maintaining its critical structural and relational integrity. (2) attack-chain causal inference: a causal inference technique is applied to identify potential attack-related edges and nodes through various weighting schemes, enhancing the graph’s interpretability and analytical value. (3) interphase attack association: a solution is implemented to assess potential relationships between partitioned subgraphs, ensuring that multi-staged attacks are effectively correlated.
IV-A2 Command-Line Differentiation Evaluation Based on Known Facts
Following the attack provenance construction phase, the system transitions to the evaluation of command-line actions. This phase integrates multiple evaluation strategies. Initially, process chains and their associated command-line executions are analyzed using Sigma rules [24]. These rules, crafted by security experts and organizations, identify specific malicious signatures, significantly reducing false positives when malicious actions match these signatures. Next, undetected command-line actions are cross-referenced against a repository of defense evasion rules [25, 26], which address common system command-line patterns often exploited for evasion. The evaluation results are then used to adjust the weights within the attack provenance graph, improving the detection accuracy.
2
Attack-Evidence Awareness. The attack-evidence awareness module is detailed below, highlighting the core elements that enable effective identification and contextualization of attack-related evidence.
IV-A3 Explainable Anomaly Detection for Unknown Attacks
As attack clauses accumulate, the corresponding evidence becomes increasingly explicit and compelling, necessitating timely detection. In the explainable anomaly detection phase, the system conducts a thorough analysis of the partitioned attack provenance graph. This process begins with Information Path (InfoPath) traversal, which identifies and ranks anomalous paths based on weighted scoring. While effective for known threats, rule-based detection alone is insufficient to capture previously unseen attack behaviors. To address this limitation, we propose a SimHash-based similarity framework that analyzes unknown command-line instances. This framework informs the selection of embedding strategies that maximize intra-class separability, thereby improving contextual awareness of novel attack patterns. To further enhance robustness, we incorporate an unsupervised voting mechanism that aggregates outputs from multiple machine learning models, reducing individual model bias and improving detection stability. This integrated approach improves anomaly detection performance while maintaining interpretability. Additionally, adaptive hyperparameter tuning is employed to optimize system parameters in real time, ensuring consistent performance across diverse and dynamic operational contexts.
IV-A4 Report Generation and Recommendation
To enhance SecDevOps efficiency, we employ the RAG-GPT4o model to automate the generation of comprehensive analysis reports derived from anomaly detection outputs. This model systematically ranks snapshot-based recommendations, highlighting the Top-5 most suspicious activities with contextual insights. These actionable insights include precise interpretations of suspicious command-line activities, inferred adversarial tactics, and potential threat indicators. By summarizing critical anomalies concisely and contextually, security analysts can quickly evaluate whether a host is compromised, significantly reducing cognitive overload and mitigating alarm fatigue. This streamlined approach empowers teams to swiftly prioritize responses, enhancing overall threat management capabilities
Figure 5: Comparison of (a) Information-Flow Focused Structure and (b) Refined Command-Line Focused Structure: Both network and command-line details are stored as memory attributes within process nodes. The refined structure omits file nodes to reduce redundancy and lower false alarm rates, as files are rarely the root cause [27].
IV-B Attack-Clause Sketch
The attack-clause module comprises two key components: (1) attack provenance construction and (2) command-line differentiation evaluation based on known attack patterns. These phases are critical for correlating attack clauses and quantifying the associated risk levels.
1
Attack Provenance Construction. To address the limitations of detecting malicious command-line activities, we design a refined attack provenance structure. This enhanced representation treats command-line inputs not merely as deviations within the global behavior graph but also as standalone semantic features. By incorporating command-line semantics as independent signals, the structure improves the granularity of behavioral differentiation and enhances the detection of subtle or previously unseen attacks. Figure 5 provides a comparative visual representation of the improvements in constructing the provenance graph.
IV-B1 Attack Provenance Structure
Figure 5(a) illustrates the conventional approach, termed the Information-Flow Focused Structure, as described in [28]. This method employs a heterogeneous graph composed of diverse node types, including reverse shell processes, system processes, command-line inputs, file paths, and network activities.
In contrast, Figure 5(b) depicts our proposed Refined Command-Line Focused Structure. Unlike the heterogeneous architecture, our design adopts an isomorphic graph structure with uniform node types, centering on system processes. Command-line and network information are encapsulated as memory attributes within each process node, reducing structural noise while preserving semantic context.
This refined design yields multiple advantages. By unifying node types, it streamlines graph representation and computation, enabling more effective application of graph-based analytics such as embedding techniques, anomaly detection algorithms, and structural similarity metrics. The reduced complexity and consistent node attributes enhance the system’s ability to detect and differentiate subtle anomalies in attack behavior.
IV-B2 Attack Irrelevance Reduction
To improve the clarity and relevance of attack provenance analysis, we introduce a reduction technique that eliminates non-correlated events from the provenance graph. This targeted reduction is applied at both the node and edge levels, mitigating the dependency explosion problem, a well-known challenge in provenance analysis [29].
At the node level, we remove subgraphs that do not contain any command-line elements. This decision is based on two key observations. First, normal system behavior seldom involves extensive command-line invocation. Second, effective incident investigation often requires the presence of command-line clauses to trace adversarial intent. By isolating command-line-relevant components, the proposed reduction method streamlines the graph, reduces analytical overhead, and minimizes the likelihood of misleading dependencies.
At the edge level, a critical concern is the presence of loops and cycles within the provenance graph, which can compromise structural integrity and impede accurate data correlation. To address this issue, we transform the graph into a Directed Acyclic Graph (DAG) [30]. This transformation involves systematically identifying and eliminating cyclic dependencies by removing looped edges, thereby ensuring acyclic structure and facilitating more reliable provenance analysis.
@@ IV-B2 Attack Irrelevance Reduction
Figure 6: Attack-chain causal inference: the process begins by calculating node weights using PageRank and Shortest-Path Betweenness Centrality. Subsequently, these node weights are used to determine the edge weights.

### v2

@@ Abstract
Endpoint Detection and Response (EDR) solutions embrace the method of attack provenance graph to discover unknown threats through system event correlation. However, this method still faces some unsolved problems in the fields of interoperability, reliability, flexibility, and practicability to deliver actionable results. Our research highlights the limitations of current solutions in detecting obfuscation, correlating attacks, identifying low-frequency events, and ensuring robust context awareness in relation to command-line activities.
To address these challenges, we introduce DefendCLI, an innovative system leveraging provenance graphs that, for the first time, delves into command-line-level detection. By offering finer detection granularity, it addresses a gap in modern EDR systems that has been overlooked in previous research. Our solution improves the precision of the information representation by evaluating differentiation across three levels: unusual system process calls, suspicious command-line executions, and infrequent external network connections. This multi-level approach enables EDR systems to be more reliable in complex and dynamic environments.
Our evaluation demonstrates that DefendCLI improves precision by approximately $1.6\times$ compared to the state-of-the-art methods on the DARPA Engagement Series attack datasets. Extensive real-time industrial testing across various attack scenarios further validates its practical effectiveness. The results indicate that DefendCLI not only detects previously unknown attack instances, which are missed by other modern commercial solutions, but also achieves a $2.3\times$ improvement in precision over the state-of-the-art research work.
@@ I Introduction
To provide a more automated security orchestration scheme, we integrate Large Language Models (LLMs) into our threat recommendation system to generate rapid preliminary assessments, enabling efficient triage before comprehensive threat analysis. Extensive experiments and evaluations show that DEFENDCLI effectively manages real-world cyber threats by identifying previously undetected attack instances.
Our contributions are summarized as follows:
@@ II Related Work
This section provides a comparative analysis of related work across four key dimensions critical for command-line activity detection: obfuscation detection (OD), attack correlation (AC), low-frequency detection (LFD), and context awareness (CA) [23]. Table I presents a feature analysis of DefendCLI relative to prior methods based on their algorithmic designs (AD).
1
OD: Challenges in Obfuscation Detection [6]. Most existing approaches primarily leverage structural anomalies, a strategy that may not fully capture the semantics of obfuscated commands (e.g., polymorphic commands or encoded payloads). Holmes [22] provides effective detection via TTP-based rules. Methods like Unicorn [7], DepComm [35], ProGrapher [36], Nodlink [17], and DISTDET [5] focus largely on structural changes or statistical deviations. While effective for general anomalies, these systems face challenges in parsing sophisticated obfuscation techniques without semantic analysis.
2
AC: Granularity in Attack Correlation [37]. Effective correlation is essential for reconstructing attack chains. Nodlink [17] demonstrates strong performance by employing Steiner trees for node-level correlation, while Holmes [22] utilizes TTP rules to provide a cohesive view of adversarial behavior. Unicorn, DepComm, and ProGrapher rely on dependency tracking, which captures structural relationships. DISTDET [5] emphasizes efficient log reduction and deduplication, which prioritizes scalability but may limit the resolution of complex relationships in multi-stage APTs.
3
LFD: Sensitivity to Low-Frequency Events [24]. Detecting rare, non-repetitive patterns is vital for identifying novel or targeted attacks. Many current methods (Unicorn, DepComm, ProGrapher, Nodlink, DISTDET) utilize historical frequency profiles, which are effective for repetitive behaviors but may be less sensitive to rare events lacking precedence. Holmes [22] offers detection through predefined rules, whereas anomaly-based baselines in other systems must balance sensitivity with false-positive rates when encountering stealthy, low-frequency threats.
4
CA: Context Awareness Integration [18]. Distinguishing benign from malicious commands often depends on execution context (e.g., user privileges, parent processes). DepComm [35] incorporates system attributes as dependencies to enhance context. Holmes [22] uses rule-based context for known attacks. Approaches such as Unicorn, ProGrapher, Nodlink, and DISTDET primarily focus on graph structure. Integrating robust environmental variables remains a key area for development to accurately differentiate between legitimate operations and malicious mimicry.
TABLE I: Feature Analysis of DEFENDCLI Relative to Prior Work. : Comprehensive, : Substantial, : Partial, : Limited.
@@
III DefendCLI
@@
The architecture of DefendCLI, illustrated in Figure 3, adopts a highly modular design to ensure extensibility and clarity in addressing the challenges of command-line attack detection. The system is fundamentally composed of two sequential computational phases: Attack-Clause Sketch, dedicated to the construction, refinement, and structural quantification of the provenance graph; and Attack-Evidence Awareness, responsible for semantic analysis, anomaly detection, and explainable threat reporting.
III-A Attack-Clause Sketch: Graph Construction and Quantification
This module focuses on transforming raw system telemetry into a structured, quantifiable representation suitable for causal analysis. It involves graph refinement, probabilistic node scoring, and interphase community association.
Figure 4: Comparison of (a) Information-Flow Focused Structure (heterogeneous nodes) and (b) Refined Command-Line Focused Structure (isomorphic process nodes with attributes). The refined structure reduces graph noise by encapsulating command and network data within process nodes.
@@ III-A Attack-Clause Sketch: Graph Construction and Quantification
Figure 5: Attack-chain causal inference: Node importance is calculated using PageRank (for rarity) and Shortest-Path Betweenness Centrality (for structural bridges). These scores are aggregated to determine initial edge weights.